OpenStudy (anonymous):

8x+9=x^2 8x+9=x^2 @Mathematics

OpenStudy (anonymous):

x = 9

OpenStudy (anonymous):

How did you get that? :/

OpenStudy (anonymous):

8x+9=x^2 x^2-8x-9=0 (x-9)(x+1)=0 x=9 and -1

OpenStudy (anonymous):

Quadratic equation can be formed, then we can find the value of x.